OpenStudy (anonymous):

Multiply. (A – 1/2) (A + 2/3)

OpenStudy (unklerhaukus):

\[=(A – \frac{1}{2})( A + \frac{2}{3})\]\[=A^2+\frac{2}{3}A-\frac{1}{2}A-\frac{1}{3}\]\[=A^2+\frac{1}{6}A-\frac{1}{3} \]

OpenStudy (anonymous):

multiply each term by the other one. for e.g. A x A, A x 2/3, -1/2 x A and -1/2 x 3. Then you collect like terms. @UnkleRhaukus shows the working out.

OpenStudy (anonymous):

thanks!